You have a total of 25 coins, all nickels ($0.05) and quarters ($0.25). The total value is $3.85. Write and solve a system of equations to find the number of nickels, n, and quarters, q
Answer by stanbon(75887)

You have a total of 25 coins, all nickels ($0.05) and quarters ($0.25). The total value is $3.85. Write and solve a system of equations to find the number of nickels, n, and quarters, q
------------------
Equations:
n + q = 25 coins
5n+25q = 385 cents
------------------------
Modify for elimination::
n + q = 25
n + 5q = 77
-----
Subtract and solve for "q"::
4q = 52
q = 13 (# of quarters)
-----
n = 25-q
n = 12 (# of nickels)
------
Cheers,
Stan H.
---------